Suppose that the economy initially has k700 in reserves.(all answers to be in kwacha at $1=k18.00)
a) If the required reserve ratio is 20%, what is the total money supply, assuming there is no cash being held? What is the money multiplier? What is the amount of loans outstanding in the banking system?
b) How much does the money supply increase when the central bank adds $5.55 in new reserve?
c) If the central bank wants to the make the money multiplier 10, what do they need to set the required reserve ratio to? If they make this change what happens to the money supply?
d) If a bank in this economy has $200 of reserves, $250 of loans and $450 of deposits, how much excess reserve are they holding (use the required reserve ratio of 10%)? How much could the bank make in additional loans?
1
Expert's answer
2021-09-30T12:21:27-0400

a)

Reserve ratio is 20%

Total reserves are K700

Reserve ratio is the percentage of total deposits that have to be kept with the central bank

"Reserve\space ratio=\frac{total\space reserves}{deposits}"


"0.2=\frac{K700}{deposits}"


Deposits=3500

Money supply=cash + deposits

"=0+K3500=K3500"

"Money\space multiplier=\frac{1}{reserve\space ratio}=\frac{1}{0.2}=5"



"Loan=deposits-rsserve \space ratio"


"=3500-700=K2800"


b)

If central bank increases $5.55 in reserves, deposits/money supply will increase by the multiplier times of $5.55

Increase in money supply will be

"5\times \$5.55=27.75"


c)

"Money\space multiplier=\frac{1}{reserve\space ratio}"


If money multiplier needs to be 10, the reserve ratio has to be 10%

So, if reserves are k700, then money supply becomes 700 * 10= 7000.


Because R70/ of reserves support R7000 of deposits at a given reserve ratio (10%)


d)

If reserves are $200 at 10% reserve ratio, it can support $2000 of deposits

$2000-$209=$1800 loans can be given.

Available loan is just $250 so additional $ 1550 can be given